You need a 80% alcohol solution. On hand, you have a 320 mL of a 85% alcohol mixture. How much pure water will you need to add to obtain the desired solution?
You will need
mL of pure water
to obtain
mL of the desired 80% solution.

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

You need a 80% alcohol solution. On hand, you have a 320 mL of a 85% alcohol mixture. How much pure water will you need to add to obtain the desired solution?
----------
Equation:
alcohol + alcohol = alcohol
0.85*320 + 0*x = 0.80(320+x)
----
Multiply thru by 100 to get:
85*320 + 0 = 80*320 + 80x
5*320 = 80x
x = 20 mL (amt. of pure water needed)
------------------------------------------
Cheers,
Stan H.
============




Answer by richwmiller(17219)

from the view of water
320*.15+w=(320+w)*.20
w=20 ml of water
from the view of alcohol
.85*320=.80*(320+w)
w=20 ml of water